OpenStudy (anonymous):

how are tangent, cotangent, secant, and cosecant related to cosine and sine?

OpenStudy (nincompoop):

DO YOU KNOW WHAT THOSE ARE?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Sin=y Cos=x Tan=y/x and csc opposite of sin. Sec opposite of cos and Cot opposite of tan.

OpenStudy (nincompoop):

ARE YOU SEEING ANY PATTERN?

OpenStudy (anonymous):

Yes but I don't know how to put it in words.

OpenStudy (nincompoop):

THEN USE SYMBOLS

OpenStudy (anonymous):

ok

OpenStudy (nincompoop):

DEFINITION SINE = Y, COSINE = X CONDITION TAN = Y/X THEN TAN = SINE/COSINE ET CETERA BLAH BLAH BLAH

OpenStudy (anonymous):

ok thank you